OpenStudy (emogirlatemoooocow):

Which expression is equivalent to - 3 x (4/- 5)? A. (- 3 x 4) / (- 3 x (-5)) B. (- 3 x 4) / 5 C. - 3/5 x 4 D. - 3/- 5 x 4

OpenStudy (misty1212):

interesting that they are not asking you to compute isn't it?

OpenStudy (emogirlatemoooocow):

yeah

OpenStudy (misty1212):

i think they want you to say \[-3\times \frac{4}{-5}=\frac{-3}{-5}\times 4\] if that is an answer choice

OpenStudy (emogirlatemoooocow):

yep its A.